i need the formula for this problem i need to fill a number of containers with a mix that is 3 cubic feet of a to 1 cubic foot b. the containers are .668 cubic feet in volume. i need to know how much of each thing i need to buy to fill 247 containers. the items only come in 2.5 cubic foot amounts for part a and 1 cubic foot of part b. i understand the total volume but how to i get from total volumes to how much of each thing?
Answer by ankor@dixie-net.com(22740)

i need the formula for this problem i need to fill a number of containers with
a mix that is 3 cubic feet of a to 1 cubic foot b. the containers are .668
cubic feet in volume.
i need to know how much of each thing i need to buy to fill 247 containers.
the items only come in 2.5 cubic foot amounts for part a and 1 cubic foot of part b.
i understand the total volume but how to i get from total volumes to how much of each thing?
:
Find the total volume required: .668 * 247 ~ 165 cu/ft
:
Let x = the multiplier, then
1x = amt of b
3x = amt of a
:
1x + 3x = 165
4x = 165
x = 165/4
x = 41.25 is multiplier
then
Mixture b = 41.25 cu/ft required
Mixture a = 3(41.25) = 123.75 cu/ft required
:
Find no. of containers of each
b = 41.25 ~ 42 containers
and
a = 123.75/2.5 = 49.5 ~ 50 containers